If 3sinx + 4cosx = 5, then 4sinx - 3cosx =....???



  1. 0

  2. -5

  3. 1/5

  4. 5


plz explain me in detail... how to do questions like this... 

Grade:12

9 Answers

KARAN ACHARYA
12 years ago

ANS    3sinx+4cosx=5

     => 3/5 sinx +4/5 cosx = 1

     let cosA=3/5   => sinA=4/5

     =>  cosAsinx + sinAcosx = 1

     =>  sin(x+A) = 1

Now,

       4sinx - 3cosx

  =   5(4/5sinx - 3/5 cosx)     [multipying numerator and denominator by 5]

  =   5(sinAsinx - cosAcosx)

  =   -5{cos(x+A)} = -5[root{1-(sin(x+A)^2)}] = -5 x 0 = 0          Ans 0

 

TRICK: if asinx+bcosx is given then multiply numerator and denominator by root(a^2 +b^2)

            this method is useful  in many questions 

 

 

 PLEASE       APPROVE       MY       ANSWER       IF       YOU       LIKE       IT

 

BARATH P
11 years ago

3sinx+4cosx=5

Squaring on both sides.

(3sinx+4cosx) 2 = 25

9sin2x +16 cos2x+24sinxcosx=25

9(1- cos2x)+16(1- sin2x)+ 24sinxcosx=25

9-9 cos2x+16-16 sin2x+24sinxcosx=25

25-9 cos2x-16 sin2x+24sinxcosx=25

9 cos2x+16 sin2x-24sinxcosx=0

(3cosx-4sinx) 2 = 0

4sinx-3cosx=0

 

Chandrabhushan Reddy Chigarapalli
8 years ago
3sinx+4cosx=5.
Do sqaring on both sides.
The equation will be as :- 9sin2x+16cos2x+24sinx.cosx=25.
Let the above equation be no.1
Think that the value of 4sinx-3cosx=p.
Do squaring on both sides.
The eqation will be as :- 16sin2x+9cos2x-24sinx.cosx=p2.
Let the above equation be no.2
Add equations no.1 and no.2
You will get as :- 25sin2x+25cos2x=25+p2
25(sin2x+cos2x)=25+p2
25=25+p2                                   (because sin2x+cos2x=1)
p2=0
p=0.
4sinx-3cosx=0.
Therefore answer is 0.
 
 
PLEASE  APPROVE  MY  ANSWER  IF IT IS RIGHT
sumanth
8 years ago
 

3sinx+4cosx=5

 

Squaring on both sides.

 

(3sinx+4cosx) 2 = 25

 

9sin2x +16 cos2x+24sinxcosx=25

 

9(1- cos2x)+16(1- sin2x)+ 24sinxcosx=25

 

9-9 cos2x+16-16 sin2x+24sinxcosx=25

 

25-9 cos2x-16 sin2x+24sinxcosx=25

 

9 cos2x+16 sin2x-24sinxcosx=0

 

(3cosx-4sinx) 2 = 0

 

4sinx-3cosx=0

Lab Bhattacharjee
8 years ago
HINT:  (3sinx +4cosx)^2+(3cosx-4sinx)^2=(3^2+4^2)(cos^2x+sin^2x)
ritesh
7 years ago
easiest method....given 3sinx+4cosx=5..then find 4sinx-3cosx=?
a=3  b=4  c=5...let our ans to be k
a^2+b^2=c^2+k^2
after putting values nd solving we get...k=0...easy method...
 
 
Kushagra Madhukar
3 years ago
Hello student
This question can be solved using two methods both are quite easy
 
Method-1
Given, 
3sinx + 4cosx = 5
Squaring both sides,
9sin2x + 16cos2x + 24sinxcosx = 25
9sin2x + 16 cos2x + 24sinxcosx = 25(sin2x + cos2x)       [ 25 = 25 x1 = 25 x (sin2x + cos2x)
24sinxcosx = 16sin2x + 9 cos2x
or, (4sinx)2 + (3cosx)2 – 2(4sinx)(3cosx) = 0
or, [4sinx – 3cosx]2 = 0
hence, 4sinx – 3cosx =0
 
Method-2
3sinx + 4cosx = 5
as we can notice- 3, 4, 5 are pythagorean triplet therefore are a part of the same right angled triangle.
Dividing both sides by 5
3/5 sinx + 4/5 cosx = 1
If we take 3/5 = cosA ; then sinA = 4/5
or, sinx cosA + cosx sinA = 1
or, sin(x + A) = 1
 
therefore,
cos(x + A) = [1 – {sin(x + A)}2]1/2  = [ 1 – 1 ]½  = 0
cosx cosA – sinx sinA = 0
or, 3/5 cosx – 4/5 sinx = 0
multiplying both sides by -5, we get,
4sinx – 3cosx = 0
 
You can use either of these methods in questions like this. But the 2nd method might be easier in case of some problems.
Hope it helps.
Regards,
Kushagra
Vikas TU
3 years ago
Dear student 
Question is not clear 
Please upload an image.
We will happy to help you.
Good luck 
Cheers 
 
Zaid
2 years ago
3sinx + 4cosx = 5
as we can notice- 3, 4, 5 are pythagorean triplet therefore are a part of the same right angled triangle.
Dividing both sides by 5
3/5 sinx + 4/5 cosx = 1
If we take 3/5 = cosA ; then sinA = 4/5
or, sinx cosA + cosx sinA = 1
or, sin(x + A) = 1
 
therefore,
cos(x + A) = [1 – {sin(x + A)}2]1/2  = [ 1 – 1 ]½  = 0
cosx cosA – sinx sinA = 0
or, 3/5 cosx – 4/5 sinx = 0
multiplying both sides by -5, we get,
4sinx – 3cosx = 0